# Effect of interatomic repulsion and quasi-degenerate states on a Kitaev-transmon qubit based on double quantum dots

### v3

@@ Acknowledgments
Appendix A Double degeneracy of the eigenstates for $E_{oo}=E_{ee}$ and $\phi_{ext}-0$
@@
$2\hat{O}\cos(\hat{\theta})|oon\rangle$ $=$ $\hat{O}(|een+1\rangle+|een-1\rangle)$
$=$ $-(-1)^{n}(|oon+1\rangle+|oon-1\rangle),$
$2\cos(\hat{\theta})\hat{O}|oon\rangle$ $=$ $-2\cos(\hat{\theta})(-1)^{n}|een\rangle$ (16)
$=$ $-(-1)^{n}(|oon+1\rangle+|oon-1\rangle).$
A similar result is obtained appying the oparators to $|een\rangle$. Therefore $\hat{O},\cos(\hat{\theta})]=0$. The remaing terms of the Hamiltonian for $E_{oo}=E_{ee}$ consist of a sum of terms of the form $H_{ij}=|eei\rangle\langle eej|+|ooi\rangle\langle eej|$. After a simple algebra one obtains

### v4

@@ Acknowledgments
Appendix A Double degeneracy of the eigenstates for $E_{oo}=E_{ee}$ and $\phi_{ext}=0$
@@
$2\hat{O}\cos(\hat{\theta})|oon\rangle$ $=$ $\hat{O}(|een+1\rangle+|een-1\rangle)=-(-1)^{n}(|oon+1\rangle+|oon-1\rangle),$
$2\cos(\hat{\theta})\hat{O}|oon\rangle$ $=$ $-2\cos(\hat{\theta})(-1)^{n}|een\rangle=-(-1)^{n}(|oon+1\rangle+|oon-1\rangle).$ (16)
A similar result is obtained appying the oparators to $|een\rangle$. Therefore $\hat{O},\cos(\hat{\theta})]=0$. The remaining terms of the Hamiltonian for $E_{oo}=E_{ee}$ consist of a sum of terms of the form $H_{ij}=|eei\rangle\langle eej|+|ooi\rangle\langle eej|$. After a simple algebra one obtains